# Signature Isolation Forest

### v1

@@ 1 Introduction
If one seeks to avoid this projection methods that take functional data to a parameter vector space to undertake the AD method, one may alternatively seek AD methods working directly in the function space. This is the case of functional data depth, which takes as input a curve and a dataset of curves and returns a score indicating how deep the curve w.r.t. the dataset is. Many functional data depths, with different formulations, have been designed so far, such as, among others, the functional Tukey depth (Fraiman and Muniz,, 2001; Claeskens et al.,, 2014), the Functional Stahel-Donoho Outlyingness (Hubert et al.,, 2015), the ACH depth (Staerman et al.,, 2020), the band depth (López-Pintado and Romo,, 2009) or the half-region depth (López-Pintado and Romo,, 2011); see Nieto-Reyes and Battey, (2016); Gijbels and Nagy, (2017) or the Chapter 3 of Staerman, (2022) for a detailed review of functional depths.
@@ 2 Background & Preliminaries
[…]
In practice, only a finite-dimensional marginal, $(X_{t_{1}},X_{t_{2}},\dots,X_{t_{p}})$ with $t_{1}<\dots<t_{p},p\geq 1$ and $(t_{1},\dots,t_{p})\in[0,1]^{p}$, can be observed. To consider a function, rather than a set of discrete values, approximation or interpolation procedures are usually used and combined with a preliminary smoothing step when the observations are noisy. Denote by $\mathcal{X}$ a reconstructed dataset of interpolated curves/functions $\mathcal{X}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\}$ from the set of observations $\{x_{i,t_{1}},x_{i,t_{2}},\dots,x_{i,t_{p}}\}_{i=1}^{n}$ (Ramsay and Silverman,, 2005). The problem of functional anomaly detection can be formulated as learning a score function $s:\mathcal{F}^{d}\rightarrow\mathbb{R}$ from $\mathcal{X}$ reflecting the degree of abnormality of any element of an infinite dimensional space $\mathcal{F}^{d}$ w.r.t. $\mathcal{X}$.
@@ 2.2 The Signature Method
$S^{k}(\mathbf{X})=\big{(}$ $1,S_{1}(\mathbf{X}),\ldots,S_{d}(\mathbf{X}),$
$S_{(1,1)}(\mathbf{X}),S_{(1,2)}(\mathbf{X}),\ldots,S_{(d,d)}(\mathbf{X}),\ldots$
$S_{(1,\ldots,1)}(\mathbf{X}),\ldots,S_{\underbrace{(d,\ldots,d)}_{k}}(\mathbf{X})\big{)}\in\mathbb{R}^{C}.$
@@
One may refer to Lee and Oberhauser, (2023) for an account of the untruncated kernel signature. See also Section A in the Appendix for further details about the signature and its properties.
@@ 3 Signature Isolation Forest Methods
A si-tree is also built by iterating this procedure until all training data curves are isolated. The two algorithms of K-SIF and SIF are summarized in the Appendix B.
@@ 4.1 Parameters Sensitivity Analysis
We investigate the behavior of K-SIF and SIF with respect to their two main parameters: the depth of the signature $k$ and the number of split windows $\omega$. For the sake of place, the experiment on the depth is postponed in Section C.1 in the Appendix.
@@
The first dataset is constructed as follows. We simulate 100 constant functions. We then select at random 90% of these curves and Gaussian noise on a sub-interval; for the remaining 10% of the curves, we add Gaussian noise on another sub-interval, different from the first one. More precisely:
  • 90% of the curves, considered as normal, are generated according to
    $\mathbf{x}(t)=b(t)+\varepsilon(t)\mathbb{I}(t\in[0.3,0.6]),$
    with $\varepsilon(t)\sim\mathcal{N}(0,1)$, $b\sim\mathcal{U}([0,100])$ and $\mathcal{U}$ representing the uniform distribution.
  • 10% of the curves, considered as abnormal, are generated according to
    $\mathbf{x}(t)=b+\varepsilon(t)\mathbb{I}(t\in[0.7,0.8]),$
    where $\varepsilon(t)\sim\mathcal{N}(0,1)$ and $b\sim\mathcal{U}([0,100])$.
The second dataset is constructed as follows. We generate 100 one-dimensional Brownian motion paths. We remark that a stochastic process $B_{t}$ follows a Brownian motion if it satisfies the following stochastic differential equation
$dB_{t}=\mu dt+\sigma dX_{t},$
where $X_{t}$ is a Wiener process or Brownian motion (with distribution $\mathcal{N}(0,1)$).
@@
Refer to caption Refer to caption
Figure 1: AUC under the ROC curve w.r.t. the number of split window on the first (left) and the second (right) datasets for the three dictionaries.
4.2 Advantages of (K-)SIF over FIF
This part highlights the advantages of our class of algorithms over Functional Isolation Forest. We provide numerical experiments supporting our approach’s appealing properties, such as the robustness to noisy data and detecting order variation changes.
Robustness to Noise. We now explore the sensitivity of our class of algorithms to noisy data. We also provide a comparison with Functional Isolation Forest. To that end, we simulate a dataset of 500 ‘relatively smooth’ standard Brownian motion paths with $\mu=0$ and $\sigma=0.05$. We also simulate 50 standard Brownian motion paths with $\mu=0$ and $\sigma\in[0.05,0.2]$. Depending on the level of noise present in the last 50 curves (i.e., standard deviation here), these paths may represent normal, noisy and abnormal data. For example, $0.05$ (corresponding to 00 noise level) is normal data, while $0.2$ (corresponding to $0.15$ noise level) is abnormal.
We perform K-SIF and FIF with both ‘Brownian’ and ‘Cosine’ dictionaries and report the AUC under the ROC curves by considering the 50 curves as the second class in Figure 2.
An AUC of 1 means that the 50 curves of the second class are identified as anomalies, while an AUC of 0.5 indicates that the algorithm cannot distinguish them from normal data. According to the findings provided in Section 4.1, the parameters for K-SIF have been chosen such that the number of split window sizes is equal to 10 while the depth of the signature $k=2$. We can observe that FIF, regardless of the dictionary used, is highly sensitive to noisy data and considers it abnormal, even with a minimal noise level. In contrast, K-SIF is way more robust to such data and requires higher noise levels to start viewing this data as anomalies.
Figure 2: Robustness of K-SIF and FIF to noise level.
(K-)SIF detects Swap Order Variation Changes. The signature method considers the order of occurring events in functional data. To investigate this phenomenon with our proposed class of algorithms, we define a synthetic dataset of $100$ smooth functions given by
$\mathbf{x}(t)=30t^{q}(1-t)^{q},$
with $t\in[0,1]$ and $q$ equispaced in $[1,1.4]$. Then, we simulate the occurrences of events by adding Gaussian noise on different portions of the functions. We randomly select 90% of them and add Gaussian values on a sub-interval, i.e.,
$\mathbf{x}(t)=30t^{q}(1-t)^{q}+\varepsilon(t)\mathbb{I}(t\in[0.2,0.4]),$
where $\varepsilon(t)\sim\mathcal{N}(0,0.8)$. We consider the 10% remaining as abnormal by adding the same ‘events’ on another sub-interval compared to the first one, i.e.,
$\mathbf{x}(t)=30t^{q}(1-t)^{q}+\varepsilon(t)\mathbb{I}(t\in[0.6,0.8]),$
where $\varepsilon(t)\sim\mathcal{N}(0,0.8)$. See Figure 8 in the Appendix for a dataset visualization. To summarize, we have constructed two identical events occurring at different parts of the functions, leading to isolating anomalies. As presented in the introduction, this class of anomalous data are amongst the most challenging to identify.
@@ 4.2 Advantages of (K-)SIF over FIF
Table 1: AUC of different anomaly detection methods calculated on the test set. Bold numbers correspond to the best result.
[…]
@@ 4.3 Real-data Anomaly Detection Benchmark
To evaluate the effectiveness of the proposed (K-)SIF algorithms and provide a comparison with FIF, we perform a comparative analysis using ten anomaly detection datasets constructed in Staerman et al., (2019) and sourced from the UCR repository (Chen et al.,, 2015). In contrast to Staerman et al., (2019), we do not use a training/test part since the labels are not used for the training and train and evaluate models on the training data only. We evaluate the algorithms’ performance by quantifying the AUC under the ROC curves.
K-SIF is examined in the context of three introduced dictionaries: Brownian, Cosine and Gaussian wavelets, as done for FIF. The parameters are configured with $N=100$, $m=\text{min}(256,n)$, the height limit set to $\lceil\log_{2}(m)\rceil$ for both FIF and K-SIF/SIF. The number of split windows is fixed with $\omega=10$, and the depth is fixed with $k=3$. We compare the K-SIF/SIF methods against two widely used multivariate anomaly detection techniques and two functional depths with default settings. The multivariate methods, namely, isolation forest (IF; Liu et al.,, 2008) and one-class support vector machine (OCSVM; Schölkopf et al.,, 2001), are applied following dimension reduction via Functional PCA. We retain 20 principal components with the largest eigenvalues using the Haar basis. The depths considered include the random projection halfspace depth (Cuevas et al.,, 2007) and the functional Stahel-Donoho outlyingness (Hubert et al.,, 2015).

### v2

@@ 1 Introduction
If one seeks to avoid this projection methods that take functional data to a parameter vector space to undertake the AD method, one may alternatively seek AD methods working directly in the function space. This is the case of functional data depth, which takes as input a curve and a dataset of curves and returns a score indicating how deep the curve w.r.t. the dataset is. Many functional data depths, with different formulations, have been designed so far, such as, among others, the functional Tukey depth (Fraiman and Muniz,, 2001; Claeskens et al.,, 2014), the Functional Stahel-Donoho Outlyingness (Hubert et al.,, 2015), the ACH depth Staerman et al., (2020), the band depth (López-Pintado and Romo,, 2009) or the half-region depth (López-Pintado and Romo,, 2011); see Nieto-Reyes and Battey, (2016); Gijbels and Nagy, (2017) or the Chapter 3 of Staerman, (2022) for a detailed review of functional depths.
@@ 2 Background & Preliminaries
Consider the functional random variable (r.v.) $\mathbf{X}$, taking values in the functional space $\mathcal{F}^{d}(I)$ of real valued functions (e.g. the Kolmogorov space $\mathcal{C}(I)$ or the Hilbert space $L_{2}(I)$) where $\textit{I}\subset\mathbb{R}$ represents an interval, such that
$\begin{split}\mathbf{X}:\Omega&\longrightarrow\mathcal{F}^{d}(\textit{I})\\ \omega&\longmapsto\mathbf{X}(\omega)=(X_{t}(\omega))_{t\in\textit{I}}.\end{split}$
@@
In practice, only a finite-dimensional marginal, $(X_{t_{1}},X_{t_{2}},\dots,X_{t_{p}})$ with $t_{1}<\dots<t_{p}$ and $(t_{1},\dots,t_{p})\in[0,1]^{p}$, can be observed. To consider a function, rather than a set of discrete values, approximation or interpolation procedures are usually used and combined with a preliminary smoothing step when the observations are noisy. Denote by $\mathcal{X}$ a reconstructed dataset of interpolated curves/functions $\mathcal{X}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\}$ from the set of observations $\{x_{i,t_{1}},x_{i,t_{2}},\dots,x_{i,t_{p}}\}_{i=1}^{n}$.The problem of functional anomaly detection can be formulated as learning a score function $s:\mathcal{F}^{d}\rightarrow\mathbb{R}$ from $\mathcal{X}$ reflecting the degree of abnormality of any element of an infinite dimensional space $\mathcal{F}^{d}$ w.r.t. $\mathcal{X}$.
@@ 2.2 The Signature Method
$S^{k}(\mathbf{X})=\big{(}$ $1,S_{1}(\mathbf{X}),\ldots,S_{d}(\mathbf{X}),S_{(1,1)}(\mathbf{X}),S_{(1,2)}(\mathbf{X}),\ldots,$
$S_{(d,d)}(\mathbf{X}),\ldots,S_{\underbrace{(d,\ldots,d)}_{k}}(\mathbf{X})\big{)}\in\mathbb{R}^{C}.$
@@
One may refer to Lee and Oberhauser, (2023) for an account of the untruncated kernel signature. See also Section 1 in the Appendix for further details about the signature and its properties.
@@ 3 Signature Isolation Forest Methods
A si-tree is also built by iterating this procedure until all training data curves are isolated. The two algorithms of K-SIF and SIF are summarized in Section 1 in the Appendix.
@@ 4.1 Parameters Sensitivity Analysis
We investigate the behavior of K-SIF and SIF with respect to their two main parameters: the depth of the signature $k$ and the number of split windows $\omega$. For the sake of place, the experiment on the depth is postponed in Section 3.1 in the Appendix.
@@
The first dataset is constructed as follows. We simulate 100 constant functions. We then select at random 90% of these curves and Gaussian noise on a sub-interval; for the remaining 10% of the curves, we add Gaussian noise on another sub-interval, different from the first one. More precisely: 90% of the curves, considered as normal, are generated according to $\mathbf{x}(t)=b+\varepsilon(t)\mathbb{I}(t\in[0.3,0.6]),$ with $\varepsilon(t)\sim\mathcal{N}(0,1)$, $b\sim\mathcal{U}([0,100])$ and $\mathcal{U}$ representing the uniform distribution; 10% of the curves, considered as abnormal, are generated according to $\mathbf{x}(t)=b+\varepsilon(t)\mathbb{I}(t\in[0.7,0.8]),$ where $\varepsilon(t)\sim\mathcal{N}(0,1)$ and $b\sim\mathcal{U}([0,100])$.
The second dataset is constructed as follows. We generate 100 one-dimensional Brownian motion paths. We remark that a stochastic process $B_{t}$ follows a Brownian motion if it satisfies the following stochastic differential equation $dB_{t}=\mu dt+\sigma dX_{t},$ where $X_{t}$ is a Wiener process or Brownian motion (with distribution $\mathcal{N}(0,1)$).
@@
Figure 1: AUC for the ROC curve w.r.t. the number of split window on the first (top) and the second (bottom) datasets for the three dictionaries.
4.2 (K-)SIF detects Swap Order Variation Changes
The signature method considers the order of occurring events in functional data. To investigate this phenomenon with our proposed class of algorithms, we define a synthetic dataset of $100$ smooth functions given by $\mathbf{x}(t)=30t^{q}(1-t)^{q},$ with $t\in[0,1]$ and $q$ equispaced in $[1,1.4]$. Then, we simulate the occurrences of events by adding Gaussian noise on different portions of the functions. We randomly select 90% of them and add Gaussian values on a sub-interval, i.e., $\mathbf{x}(t)=30t^{q}(1-t)^{q}+\varepsilon(t)\mathbb{I}(t\in[0.2,0.4]),$ where $\varepsilon(t)\sim\mathcal{N}(0,0.8)$. We consider the 10% remaining as abnormal by adding the same ‘events’ on another sub-interval compared to the first one, i.e., $\mathbf{x}(t)=30t^{q}(1-t)^{q}+\varepsilon(t)\mathbb{I}(t\in[0.6,0.8]),$ where $\varepsilon(t)\sim\mathcal{N}(0,0.8)$. See Figure 5 in the Appendix for a dataset visualization. To summarize, we have constructed two identical events occurring at different parts of the functions, leading to isolating anomalies. As presented in the introduction, this class of anomalous data are amongst the most challenging to identify.
@@
Table 1: AUROC of different anomaly detection methods calculated on the test set. Bold numbers correspond to the best result (higher is better).
Methods : SIF K-SIFGW FIFGW K-SIFC FIFC K-SIFB FIFB IF OCSVM fHD fSDO
@@ 4.3 Real-data Anomaly Detection Benchmark
To evaluate the effectiveness of the proposed (K-)SIF algorithms and provide a comparison with FIF, we perform a comparative analysis using ten anomaly detection datasets constructed in Staerman et al., (2019) and sourced from the UCR repository (Chen et al.,, 2015). In contrast to Staerman et al., (2019), we do not use a training/test part since the labels are not used for the training and train and evaluate models on the training data only. We evaluate the algorithms’ performance by quantifying the Area under the ROC curves. For completeness, the FPR at 95% TPR and the Area under the PR curves are given in Tables 2 and 3 in the Appendix.
K-SIF is examined in the context of three introduced dictionaries: Brownian, Cosine and Gaussian wavelets, as done for FIF. The parameters are configured with $N=100$, $m=\text{min}(256,n)$, the height limit set to $\lceil\log_{2}(m)\rceil$ for both FIF and K-SIF/SIF. The number of split windows is fixed with $\omega=10$, and the depth is fixed with $k=3$. We compare the K-SIF/SIF methods against two widely used multivariate anomaly detection techniques and two functional depths with default settings. The multivariate methods, namely, Isolation Forest (IF; Liu et al.,, 2008) and one-class support vector machine (OCSVM; Schölkopf et al.,, 2001), are applied following dimension reduction via Functional PCA. We retain 20 principal components with the largest eigenvalues using the Haar basis. The depths considered include the random projection halfspace depth (Cuevas et al.,, 2007) and the functional Stahel-Donoho outlyingness (Hubert et al.,, 2015). Due to the relative small size of datasets, see Table 1 in the Appendix, deep learning based method performs poorly as shown in Tables 5, 6 and 7 in the Appendix.